I hope I put this under the correct topic.
One recipe calls for 1/4 teaspoon vanilla. A second recipe requires for 1/2 teaspoon vanilla, and a third recipe calls for 1/8 teaspoon vanilla. How much vanilla is required to make all three recipes?

Answer by checkley71(8403)

1/2+1/4+1/8=4/8+2/8+1/8=7/8 TEASPOONS